Henry drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 8

 hours. When Henry drove home, there was no traffic and the trip only took 6

 hours. If his average rate was 16

 miles per hour faster on the trip home, how far away does Henry live from the mountains?

Do not do any rounding.


1
Expert's answer
2021-05-13T11:02:13-0400

distance = rate*time ... d = rt

 

Going: d = 8r

Returning: d = 6(r+18)

 

Since the distance is the same set them equal and solve for r.

 

8r = 6(r+18)

8r = 6r + 108

2r = 108

r = 54 mph

 

d = 8r = 8(54) =432 miles

 

check: 6(r+18) = 6(54+18) = 6(72) = 432 miles